OpenStudy (anonymous):

algebra help? write a function rule for the area of a triangle with a base of 3 cm greater than 5 times its height. what is the area of a triangle when its height is 6 cm?

OpenStudy (mathmale):

Algebra! Translate "3cm greater than 5 times the height" into b=5h+3, where b is the base and h is the height. Now suppose h=the height=6 cm. Calculate the base. What is the formula for the area of a triangle when the base and height are known? woofwoof

OpenStudy (anonymous):

I'm so confused. haha, im sorry, i must be a pain... >.< meow meow c:

OpenStudy (mathmale):

"a base of 3 cm greater than 5 times its height" is practically an invitation to choose variables and then represent the relationship between base and height algebraically. This is one of the foundation stones of algebra. I use b for base and h for height. I derive a formula for b in terms of h by translating those words into symbols: a base of 3 cm greater than 5 times its height => b=5h+3. OK: if h is 6cm, b=5(6cm)+3cm= ???

OpenStudy (anonymous):

5(6cm)+3cm= 33 cm?